Hi
steve and andrew had $540 altogether. Steve gave 1/7 of his money to andrew. In return andrew gave 1/4 of the total amount he had to steve. Then they had both an equal amount of money. How much did steve have at first.
thanks

Answer by jorel1380(3719)

Let s and a be steve and andrew, respectively. Then s+a=540. If steve gives 1/7 of his money to andrew, and andrew gives 1/4 of his money to steve, leaving them with equal amounts of money. So:
(6/7)s=(3/4)a
6s/7=3a/4
24s=21a
s+a=540
24s-21a=0
21s+21a=11340
45s=11340
s=252
a=288
☺☺☺☺